14. The mean salary of 24 people is $9000. How much will one additional salary of $80,000 increase the mean salary?
Answer by Mathtut(3670)

24(9000)=216000
:
216000+80000=296000
:
296000/25=11840
:
11840-9000=2840 increase in mean salary of( 25 people)